# Membership Inference Attacks on Large-Scale Models: A Survey

### v1

@@ 7 Conclusion
With the rapid advancement of technology, large-scale machine learning models, including Large Language Models (LLMs) and multimodal models, are increasingly being deployed across various domains. In this work, we covered most existing studies about MIA against LLMs and multimodal models, systematically reviewing each study, categorizing different MIA approaches, and analyzing their strengths and limitations. Furthermore, given the widespread adoption of fine-tuning techniques in model deployment, we also examined current fine-tuning MIA against LLMs and multimodal models. Based on the research gaps identified in the survey, we proposed several directions for future research, aiming to improve the practicality, effectiveness, and robustness of MIAs in real-world scenarios. We hope that this work serves as a valuable resource for researchers and developers working in machine learning security and privacy, fostering further advancements in the field.

### v2

@@ 7 Conclusion
With the rapid advancement of technology, large-scale machine learning models, including LLMs and LMMs, are increasingly being deployed across various domains. In this work, we covered most existing studies about MIA against LLMs and LMMs, systematically reviewing each study, categorizing different MIA approaches, and analyzing their strengths and limitations. Furthermore, given the widespread adoption of fine-tuning and other pipelines in model development and deployment, we also examined current MIA against LLMs and LMMs across pipelines. Based on the research gaps identified in the survey, we proposed several directions for future research, aiming to improve the practicality, effectiveness, and robustness of MIAs in real-world scenarios. We hope that this work serves as a valuable resource for researchers and developers working in machine learning security and privacy, fostering further advancements in the field.